# Geometric foundations of thermodynamics in the quantum regime

### v1

@@ X Connections, curvature and holonomy in the quantum thermodynamic fiber bundle
The curvature is a $\mathscr{V}$-valued two-form, expressed via the covariant derivative of the connection $\nabla$: $R(X,Y)=\nabla_{X}Y-\nabla_{Y}X-[X,Y]$. In a local trivialization of the bundle, let $\mathcal{D}$ have coordinates $\{b_{k}\}_{k=1}^{n}$, and fibers have coordinates $\{v_{\alpha}\}_{\alpha=0}^{n}$, so $\mathcal{M}$ has coordinates $(b_{k},v_{\alpha})$, and $\Xi(b_{k},v_{\alpha})=\sigma(b_{k})$ The vertical subbundle is $\mathscr{V}=\mathrm{span}\{\partial_{v_{\alpha}}\}_{\alpha=0}^{n}$. Horizontal vector fields are
@@
Holonomy arises when parallel transporting along a closed loop in $\mathcal{D}$.
@@
For a closed loop $\gamma:S^{1}\to\mathcal{D}$ based at $\sigma\in\mathcal{D}$, with horizontal lift $\tilde{\gamma}:S^{1}\to\mathcal{M}$ starting at $p_{0}\in\Xi^{-1}(\sigma)$, the holonomy is the vertical displacement
@@
The idea is that for a closed loop $\gamma$ in $\mathcal{D}$, the horizontal lift starting at $p_{0}\in\Xi^{-1}(\sigma)$ may end at a different point $p_{1}\equiv\tilde{\gamma}(1)\in\Xi^{-1}(\sigma)$, with the displacement $h=p_{1}-p_{0}\in\mathscr{V}_{p_{0}}$. The holonomy is an element of the fiber structure group, acting as a translation in the thermodynamic variables.
@@
For a closed loop $\gamma$ bounding a surface $\mathscr{S}\subset\mathcal{D}$, the holonomy is
@@
where the integral is the fiber-valued integral of the curvature two-form over $\mathscr{S}$.
@@
The connection one-form $\omega\in\Lambda^{1}(\mathcal{D})$ satisfies $\nabla_{X}Y=[X,Y]-\omega(X)Y$ for horizontal vector fields $X,Y.$ The curvature is $R=d\omega+\omega\wedge\omega$. For a loop $\gamma$, the holonomy is the path-ordered exponential
@@
where $\mathcal{P}$ denotes path ordering along $\gamma$, and $\omega$ is the $\mathfrak{g}$-valued connection 1-form. For abelian structure groups (e.g. translations in $\mathbb{R}^{n+1}$), this simplifies to
@@
Even in the non-abelian case, holonomy represents a geometric phase acquired during a cyclic evolution of the intensive parameters $\boldsymbol{\lambda}$. Upon returning to the same quantum state $\sigma$, the thermodynamic labels $(S,\boldsymbol{a})$ are transformed by a non-trivial group element $h\neq e$, for all $h\in G$, and $e$ the identity element of $G$. This transformation is not removable by local reparameterization and induces irreversible entropy production when the system is forced to return to its initial thermodynamic state.
@@
Any cyclic thermodynamic process corresponding to a closed loop $\gamma$ in $\mathcal{D}$ with non-zero holonomy $\mathrm{Hol}(\gamma,p_{0})\neq 0$ induces geometric irreversibility, contributing to net entropy production $\Sigma_{\gamma}>0$, unless the connection is flat, reflecting irreversibility due to curvature.
@@
If the connection is flat $(R=0)$, cyclic processes in $\mathcal{D}$ are geometrically reversible: the horizontal lift of $\gamma$ returns to the initial point in the fiber, with zero geometric entropy production $\Sigma_{\gamma}=0$.
This means cyclic processes (loops in parameter space like temperature) may return the quantum state but shift thermodynamic variables (e.g. entropy), leading to irreversibility—requiring dissipation to correct the shift.
@@
In the abelian case, this implies that holonomy defines a group homomorphism $\Phi:\pi_{1}(\mathcal{B})\to G$, where $\Phi([\gamma])=h(\gamma,p_{0})$, independent of base point $p_{0}$, since vertical translations commute. The image of $\Phi$ is generated by curvature integrals over a basis of 2-cycles in $H_{2}(\mathcal{B})$, via de Rham cohomology.
@@
Two cyclic protocols in intensive parameter space $\boldsymbol{\lambda}(t)$ that return to the same Gibbs state $\sigma$ but traverse homotopically distinct paths accumulate different geometric entropy shifts, even in the quasistatic limit. This constitutes a topological contribution to irreversibility—unremovable by slowing down the process—and sets a fundamental lower bound on dissipation in cyclic processes operating over topologically non-trivial control manifolds. The topological irreversibility is then tied to the concept of topological entropy production.
In the current framework, where $\mathcal{B}$ is contractible in many finite-dimensional models (e.g., full-rank Gibbs states), $\pi_{1}(\mathcal{B})=0$, and holonomy vanishes for contractible loops. However, in constrained ensembles, $\pi_{1}(\mathcal{B})$ may be non-trivial, opening the door to topological quantum thermodynamics.
@@ XI.1 Principal bundle
To elevate the quantum thermodynamic fiber bundle $(\mathcal{M},\mathcal{B},\Xi,F)$ to a principal bundle, we recall the definition of a principal $G$-bundle. A principal $G$-bundle consists of a total space $P$, a base manifold $B$, a smooth surjective projection $\pi:P\to B$, and a Lie group $G$ acting smoothly, freely, and transitively on the right on each fiber $\pi^{-1}(b)\cong G$, such that $B=P/G$ is the orbit space. The bundle admits local trivializations $\pi^{-1}(U)\cong U\times G$ for an open cover $\{U_{i}\}$ of $B$, with transition functions taking values in $G$.
@@
$(S,\boldsymbol{a},\boldsymbol{\lambda})\cdot g=(S+\delta S,\boldsymbol{a}+\delta\boldsymbol{a},\boldsymbol{\lambda}),$ (70)
for $g=(\delta S,\delta\boldsymbol{a})\in\mathbb{R}^{n+1}$. This action is free, since $p\cdot g=p$ implies $g=(0,\boldsymbol{0})$, the identity. It is transitive on each fiber $F_{\sigma}=\Xi^{-1}(\sigma)$, because for any two points $p_{1}=(S_{1},\boldsymbol{a}_{1},\boldsymbol{\lambda})$, $p_{2}=(S_{2},\boldsymbol{a}_{2},\boldsymbol{\lambda})\in F_{\sigma}$, the group element $g=(S_{2}-S_{1},\boldsymbol{a}_{2}-\boldsymbol{a}_{1})$ satisfies $p_{1}\cdot g=p_{2}$. The action preserves the fibers (fixed $\boldsymbol{\lambda}$) and commutes with the projection $\Xi$, since $\Xi(S,\boldsymbol{a},\boldsymbol{\lambda})=\rho_{\boldsymbol{\mu}(S,\boldsymbol{a},\boldsymbol{\lambda})}$ depends only on $\boldsymbol{\mu}$, which is invariant under translations in $(S,\boldsymbol{a})$.
The base manifold $\mathcal{B}$ is the quotient $\mathcal{M}/\mathbb{R}^{n+1}$, with each orbit  [^1] The quotient map $\mathcal{M}\rightarrow\mathcal{M}/\mathbb{R}^{n+1}=\mathcal{B},\ p\;\mapsto\;[p]=\operatorname{Orb}(p)$ sends every point to its orbit, defined as the set of all points in $\mathcal{M}$ that are reached from a given starting point by acting with the group $\mathbb{R}^{n+1}$. Since the group acts freely and transitively on each fiber, the orbit of any point $p\in\mathcal{M}$ is exactly the whole fiber that contains $p$. (i.e., each equivalence class) corresponding to a fiber $F_{\sigma}$. Local triviality follows from the global coordinate structure $\mathcal{M}\cong\mathbb{R}^{n+1}\times\mathbb{R}^{n}$, where the first factor corresponds to the group $G$ and the second parametrizes $\mathcal{B}$. Thus, $(\mathcal{M},\mathcal{B},\Xi,\mathbb{R}^{n+1})$ is a smooth principal $\mathbb{R}^{n+1}$-bundle with abelian structure group. ∎
This principal structure interprets the redundancy in thermodynamic labeling as a gauge symmetry: group actions correspond to translations in $(S,\boldsymbol{a})$ at fixed $\boldsymbol{\lambda}$, preserving the physical quantum state $\sigma$ but altering its thermodynamic description—analogous to gauge freedom in field theories, where different potentials describe the same physical field.
@@
    This means the connection is invariant under global translations in the fiber coordinates $(S,\boldsymbol{a})$.
@@
    Since $\exp(tX)=tX$ for $G=\mathbb{R}^{n+1}$, in coordinates we have $\xi_{X}=X^{0}\partial S+\sum_{i=1}^{n}X^{i}\partial a_{i}$. The connection must satisfy
@@
The principal $\mathbb{R}^{n+1}$-bundle structure $(\mathcal{M},\mathcal{B},\Xi,\mathbb{R}^{n+1})$ reveals that the redundancy in thermodynamic labeling: multiple values of $(S,\boldsymbol{a})$ corresponding to the same physical quantum state $\sigma\in\mathcal{B}$ at fixed intensive parameters $\boldsymbol{\lambda}$ is a gauge symmetry.
@@
Let $\Xi:\mathcal{M}\to\mathcal{B}$ be the principal quantum thermodynamic bundle. A global gauge transformation is a bundle automorphism $f:\mathcal{M}\to\mathcal{M}$, i.e. a diffeomorphism satisfying:
@@
In local coordinates $(S,\boldsymbol{a},\boldsymbol{\lambda})$ on $\mathcal{M}$, a gauge transformation $f\in\mathcal{G}(\mathcal{M})$ acts as
$f(S,\boldsymbol{a},\boldsymbol{\lambda})=\bigl(S+\phi_{S}(\boldsymbol{\lambda}),\boldsymbol{a}+\boldsymbol{\phi}_{\boldsymbol{a}}(\boldsymbol{\lambda}),\boldsymbol{\lambda}\bigr),$ (74)
where $\phi_{S}:\mathcal{B}\to\mathbb{R}$ and $\boldsymbol{\phi}_{\boldsymbol{a}}:\mathcal{B}\to\mathbb{R}^{n}$ are smooth functions. This translates thermodynamic labels within each fiber while preserving the physical state: $\Xi(f(p))=\Xi(p)=\sigma$; and the intensive parameters: $\boldsymbol{\lambda}(f(p))=\boldsymbol{\lambda}(p)$.
Assuming that the Gibbs map is injective, see Remark 9 for details, the following can be stated:
Proposition 12.
The injectivity of the Gibbs map $\boldsymbol{\lambda}\mapsto\rho_{\boldsymbol{\lambda}}$ implies that the equilibrium submanifold $\mathcal{E}$ intersects each fiber $F_{\sigma}$ at exactly one point. Then, equilibrium states are gauge-fixed.
Proof.
For $\sigma=\rho_{\boldsymbol{\lambda}}$, the equilibrium point is $p_{\sigma}=\bigl(S(\sigma),\boldsymbol{a}(\sigma),\boldsymbol{\lambda}\bigr)\in\mathcal{E}$. By injectivity, $\boldsymbol{\lambda}$ is unique. By the maximum entropy principle, $S(\sigma)$ and $\boldsymbol{a}(\sigma)$ are uniquely determined. Any other point in $F_{\sigma}$ has incorrect $(S,\boldsymbol{a})$ and lies off $\mathcal{E}$. ∎
@@
The horizontal distribution can be defined as the kernel of $\omega$:
@@
XI.2 Curvature
In a local trivialization $\Xi^{-1}(U)\cong U\times\mathbb{R}^{n+1}$ of the principal $\mathbb{R}^{n+1}$-bundle $(\mathcal{M},\mathcal{B},\Xi,\mathbb{R}^{n+1})$, let $U\subset\mathcal{B}$ be an open set with coordinates $\{\lambda_{k}\}_{k=1}^{n}$ (so $b_{k}=\lambda_{k}$), and let the fiber $\mathbb{R}^{n+1}$ have coordinates $\{v_{\alpha}\}_{\alpha=0}^{n}$, where $v_{0}=S,$ and $v_{i}=a_{i},\ \text{for }i=1,\dots,n$. The left-invariant Maurer–Cartan form $\theta$ on the abelian Lie group $G=(\mathbb{R}^{n+1},+)$ is

### v2

@@ X.1 Ehresmann connection
Example 3.
Return to the qubit example. The quantum thermodynamic state space $\mathcal{M}\cong\mathbb{R}^{3}$ has coordinates $(S,a,\lambda)$. We choose the function $\mu(S,a,\lambda)=\lambda$. The fibers $F_{\rho_{\lambda}}=\Xi^{-1}(\rho_{\lambda})$ are the level sets of $\mu$, which are 2-dimensional submanifolds diffeomorphic to $\mathbb{R}^{2}$ with coordinates $S,a$ at fixed $\lambda$. The equilibrium submanifold $\mathcal{E}$ is the 1-dimensional Legendrian curve parameterized by $\lambda$
$a(\lambda)=\tanh\lambda,\ S(\lambda)=\log(2\cosh\lambda)-\lambda\tanh\lambda.$ (58)
The contact form vanishes on $\mathcal{E}$, since
$\frac{da}{d\lambda}=\mathrm{sech}^{2}\lambda,\ \frac{dS}{d\lambda}=\lambda\,\mathrm{sech}^{2}\lambda,$ (59)
so $dS=\lambda da$, as expected.
The Bures–Wasserstein metric on $\mathcal{E}$ is the quantum Fisher information matrix. The symmetric logarithmic derivative is directly $L=\partial_{\lambda}\log\rho_{\lambda}$, which satisfies $\rho_{\lambda}L+L\rho_{\lambda}=2\partial_{\lambda}\rho_{\lambda}$, where $L=-\sigma_{z}+\tanh\lambda\ \mathrm{Id}$. The metric component is then
$g_{\lambda\lambda}(\lambda)=\operatorname{\mathrm{tr}}(\rho_{\lambda}L^{2})=1-\tanh^{2}\lambda=\mathrm{sech}^{2}\lambda.$ (60)
For a quasistatic path $\gamma(t)=(S,a,\lambda(t))$ on $\mathcal{E}$, the thermodynamic length is
$L(\gamma)=\int_{0}^{T}\sqrt{g_{\lambda\lambda}(\lambda(t))(\lambda^{\prime}(t))^{2}}dt=\int_{0}^{T}|\lambda^{\prime}(t)|\mathrm{sech}(\lambda(t))dt.$ (61)
Geodesics minimize this length and correspond to optimal quasistatic transformations.
For non-equilibrium dynamics, consider a point $p=(S_{0},a_{0},\lambda_{0})\in F_{\rho_{\lambda_{0}}}\setminus\mathcal{E}$. A relaxation path within the fiber is $\gamma(t)=(S(t),a(t),\lambda_{0})$ with, e.g., a linear relaxation dependence
$S^{\prime}(t)=-\kappa_{S}(S(t)-S(\lambda_{0})),\quad a^{\prime}(t)=-\kappa_{a}(a(t)-a(\lambda_{0})),$ (62)
$\kappa_{S},\kappa_{a}>0$, converging exponentially to the equilibrium point as $t\to\infty$. To quantify the dissipation, the metric on $\mathcal{M}$ reads
$g_{\mathcal{M}}=C^{-1}dS^{2}+\chi^{-1}da^{2}+g_{\lambda\lambda}d\lambda^{2}+\alpha\lambda(dS\otimes d\lambda+d\lambda\otimes dS),$ (63)
where $C>0$ and $\chi>0$ are capacity- and susceptibility-like functions (e.g., $C=\mathrm{sech}^{2}\lambda$, $\chi=1$), and $\alpha\in\mathbb{R}$ controls coupling. On $\mathcal{E}$, $dS=\lambda da$ ensures $g_{\mathcal{M}}|_{\mathcal{E}}=g_{\lambda\lambda}d\lambda^{2}=g_{\mathrm{BW}}$. Restricted to the fiber ($d\lambda=0$),
$g_{\mathcal{M}}|_{F_{\rho_{\lambda_{0}}}}=C^{-1}dS^{2}+\chi^{-1}da^{2}.$ (64)
The thermodynamic length of the relaxation path is
$L(\gamma)=\int_{0}^{\infty}\sqrt{C^{-1}(S^{\prime})^{2}+\chi^{-1}(a^{\prime})^{2}}dt.$ (65)
Assuming constant $C,\chi$ and $\kappa_{S}=\kappa_{a}=\kappa$,
$L(\gamma)=\sqrt{\frac{(S_{0}-S(\lambda_{0}))^{2}}{C}+\frac{(a_{0}-a(\lambda_{0}))^{2}}{\chi}}.$ (66)
The instantaneous entropy production rate is $\varsigma(t)=\kappa g_{\mathcal{M}}(\gamma^{\prime},\gamma^{\prime})$, and the entropy production reads
$\Sigma=\kappa\left(\frac{(S_{0}-S(\lambda_{0}))^{2}}{2C}+\frac{(a_{0}-a(\lambda_{0}))^{2}}{2\chi}\right)>0,$ (67)
vanishing only at equilibrium.
After relaxation, quasistatic transformations between $\lambda_{0}$ and $\lambda_{1}$ follow geodesics on $\mathcal{E}$, with length
$L=\int_{\lambda_{0}}^{\lambda_{1}}\mathrm{sech}u\ du=\arctan(\sinh\lambda_{1})-\arctan(\sinh\lambda_{0}),$ (68)
minimal and reversible (zero excess entropy production) by the properties of the Riemannian manifold $(\mathcal{B},g_{\mathrm{BW}})$.
This simple example captures the full bundle geometry: fibers encode non-equilibrium redundancy, equilibrium is the unique gauge-fixed section, quasistatic processes are horizontal geodesics, and relaxation incurs geometric dissipation bounded by the pseudo-Riemannian structure.
X.2 Curvature
@@
The curvature is a $\mathscr{V}$-valued two-form, expressed via the covariant derivative of the connection $\nabla$: $R(X,Y)=\nabla_{X}Y-\nabla_{Y}X-[X,Y]$. In a local trivialization of the bundle, let $\mathcal{B}$ have coordinates $\{b_{k}\}_{k=1}^{n}$, and fibers have coordinates $\{v_{\alpha}\}_{\alpha=0}^{n}$, so $\mathcal{M}$ has coordinates $(b_{k},v_{\alpha})$, and $\Xi(b_{k},v_{\alpha})=\sigma(b_{k})$ The vertical subbundle is $\mathscr{V}=\mathrm{span}\{\partial_{v_{\alpha}}\}_{\alpha=0}^{n}$. Horizontal vector fields are
@@
X.3 Holonomy
Holonomy arises when parallel transporting along a closed loop in $\mathcal{B}$.
@@ X.3 Holonomy
For a closed loop $\gamma:S^{1}\to\mathcal{B}$ based at $\sigma\in\mathcal{B}$, with horizontal lift $\tilde{\gamma}:S^{1}\to\mathcal{M}$ starting at $p_{0}\in\Xi^{-1}(\sigma)$, the holonomy is the vertical displacement
@@
The idea is that for a closed loop $\gamma$ in $\mathcal{B}$, the horizontal lift starting at $p_{0}\in\Xi^{-1}(\sigma)$ may end at a different point $p_{1}\equiv\tilde{\gamma}(1)\in\Xi^{-1}(\sigma)$, with the displacement $h=p_{1}-p_{0}\in\mathscr{V}_{p_{0}}$. The holonomy is an element of the fiber structure group, acting as a translation in the thermodynamic variables.
@@
For a closed loop $\gamma$ bounding a surface $\mathscr{S}\subset\mathcal{B}$, the holonomy is
@@
where the integral is the fiber-valued integral of the curvature form over $\mathscr{S}$ for the abelian case.
@@
The connection one-form $\omega\in\Lambda^{1}(\mathcal{B})$ satisfies $\nabla_{X}Y=[X,Y]-\omega(X)Y$ for horizontal vector fields $X,Y.$ The curvature is $R=d\omega+\omega\wedge\omega$, such that for a loop $\gamma$, the holonomy is the path-ordered exponential
@@
where $\mathcal{P}$ denotes path ordering along $\gamma$, and $\omega$ is the $\mathfrak{g}$-valued connection 1-form. For abelian structure groups (e.g., translations in $\mathbb{R}^{n+1}$), this simplifies to
@@
Even in the non-abelian case, holonomy represents a geometric phase acquired during a cyclic evolution of the parameters $\bar{\boldsymbol{\lambda}}$. Upon returning to the same quantum state $\sigma$, the thermodynamic labels $(\bar{S},\bar{\boldsymbol{a}})$ are transformed by a non-trivial group element $h\neq e$, for all $h\in G$, and $e$ is the identity element of $G$. This transformation is not removable by local reparameterization and induces irreversible entropy production when the system is forced to return to its initial thermodynamic state.
@@
Any cyclic thermodynamic process corresponding to a closed loop $\gamma$ in $\mathcal{B}$ with non-zero holonomy $\mathrm{Hol}(\gamma,p_{0})\neq 0$ induces geometric irreversibility, contributing to net entropy production $\Sigma\neq 0$, unless the connection is flat, reflecting irreversibility due to curvature.
@@
If the connection is flat $(R=0)$, cyclic processes in $\mathcal{B}$ are geometrically reversible: the horizontal lift of $\gamma$ returns to the initial point in the fiber, with zero geometric entropy production $\Sigma=0$.
This means cyclic processes (loops in parameter space like temperature) may return the quantum state but shift thermodynamic variables (e.g., entropy), leading to irreversibility—requiring dissipation to correct the shift.
@@
In the abelian case, this implies that holonomy defines a group homomorphism $\Phi:\pi_{1}(\mathcal{B})\to G$, where $\Phi([\gamma])=\mathrm{Hol}(\gamma,p_{0})$, independent of base point $p_{0}$, since vertical translations commute. The image of $\Phi$ is generated by curvature integrals over a basis of 2-cycles in $H_{2}(\mathcal{B})$, via de Rham cohomology.
@@
Two cyclic protocols in intensive parameter space $\bar{\boldsymbol{\lambda}}(t)$ that return to the same Gibbs state $\sigma$ but traverse homotopically distinct paths accumulate different geometric entropy shifts, even in the quasistatic limit. This constitutes a topological contribution to irreversibility—unremovable by slowing down the process—and sets a fundamental lower bound on dissipation in cyclic processes operating over topologically non-trivial control manifolds. The topological irreversibility is then tied to the concept of topological entropy production.
In the current framework, where $\mathcal{B}$ is contractible in many finite-dimensional models (e.g., full-rank Gibbs states), $\pi_{1}(\mathcal{B})=0$, and holonomy vanishes for contractible loops. However, in constrained ensembles, $\pi_{1}(\mathcal{B})$ may be non-trivial, opening the door to topological quantum thermodynamics. These topological considerations are further expanded in Sec. XI.4 by elevating the bundle to a principal bundle structure.
The vertical subbundle corresponds to shifts in thermodynamic labels at fixed quantum state, modeling internal adjustments without altering the density operator. On the other hand, the horizontal subbundle enables consistent evolution of labels as the state changes, akin to parallel transport in gauge theories. The pseudo-Riemannian metric $g_{\mathcal{M}}$ extends the Bures-Wasserstein metric off equilibrium, with a potentially indefinite signature along the entropy direction suggesting a timelike character reminiscent of spacetime metrics. Curvature measures the non-integrability of horizontal fields, sourcing holonomy as a thermodynamic geometric phase in cyclic processes, which shifts labels (e.g., inducing an extra $\Delta\bar{S}$), and necessitates dissipative resets to close the thermodynamic cycle. Flat connections permit fully reversible cycles without geometric losses, whereas non-zero curvature introduces irreducible irreversibility even in quasistatic limits. The contact non-integrability and positive-definite aspects on $\ker(\eta)$ underpin the second law, ensuring some sort of non-zero entropy production. Therefore, the unification of thermodynamics with gauge-like principles offers analogs to effects like the Aharonov-Bohm phase for electrodynamics. From a practical standpoint, the direct consequences involve geometric dissipation in cyclic processes in the quantum regime, such as quantum heat engines.
@@ XI.1 Principal bundle
To elevate the quantum thermodynamic fiber bundle $(\mathcal{M},\mathcal{B},\Xi,F)$ to a principal bundle, we recall the definition of a principal $G$-bundle.
Definition 22.
A principal $G$-bundle consists of a total space $P$, a base manifold $B$, a smooth surjective projection $\pi:P\to B$, and a Lie group $G$ acting smoothly, freely, and transitively on the right on each fiber $\pi^{-1}(b)\cong G$, such that $B=P/G$ is the orbit space. The bundle admits local trivializations $\pi^{-1}(U)\cong U\times G$ for an open cover $\{U_{i}\}$ of $B$, with transition functions taking values in $G$.
@@
$(\bar{S},\bar{\boldsymbol{a}},\bar{\boldsymbol{\lambda}})\cdot g=(\bar{S}+\delta\bar{S},\bar{\boldsymbol{a}}+\delta\bar{\boldsymbol{a}},\bar{\boldsymbol{\lambda}}),$ (80)
for $g=(\delta\bar{S},\delta\bar{\boldsymbol{a}})\in\mathbb{R}^{n+1}$. This action is free, since $p\cdot g=p$ implies $g=(0,\boldsymbol{0})$, the identity. It is transitive on each fiber $F_{\sigma}=\Xi^{-1}(\sigma)$, because for any two points $p_{1}=(\bar{S}_{1},\bar{\boldsymbol{a}}_{1},\bar{\boldsymbol{\lambda}})$, $p_{2}=(\bar{S}_{2},\bar{\boldsymbol{a}}_{2},\bar{\boldsymbol{\lambda}})\in F_{\sigma}$, the group element $g=(\bar{S}_{2}-\bar{S}_{1},\bar{\boldsymbol{a}}_{2}-\bar{\boldsymbol{a}}_{1})$ satisfies $p_{1}\cdot g=p_{2}$. The action preserves the fibers (fixed $\bar{\boldsymbol{\lambda}}$).
The base manifold $\mathcal{B}$ is the quotient $\mathcal{M}/\mathbb{R}^{n+1}$, with each orbit  [^1] The quotient map $\mathcal{M}\rightarrow\mathcal{M}/\mathbb{R}^{n+1}=\mathcal{B},\ p\;\mapsto\;[p]=\operatorname{Orb}(p)$ sends every point to its orbit, defined as the set of all points in $\mathcal{M}$ that are reached from a given starting point by acting with the group $\mathbb{R}^{n+1}$. Since the group acts freely and transitively on each fiber, the orbit of any point $p\in\mathcal{M}$ is exactly the whole fiber that contains $p$. (i.e., each equivalence class) corresponding to a fiber $F_{\sigma}$. Local triviality follows from the global coordinate structure $\mathcal{M}\cong\mathbb{R}^{n}\times\mathbb{R}^{n+1}$, where the first factor parametrizes $\mathcal{B}$ and the second corresponds to the group $G$. Thus, $(\mathcal{M},\mathcal{B},\Xi,\mathbb{R}^{n+1})$ is a smooth principal $\mathbb{R}^{n+1}$-bundle with abelian structure group. ∎
This principal structure interprets the redundancy in thermodynamic labeling as a gauge symmetry: group actions correspond to translations in $(\bar{S},\bar{\boldsymbol{a}})$ at fixed $\bar{\boldsymbol{\lambda}}$, preserving the physical quantum state $\sigma$ but altering its thermodynamic description—analogous to gauge freedom in field theories, where different potentials describe the same physical field.
@@
    This means the connection is invariant under global translations in the fiber coordinates $(\bar{S},\bar{\boldsymbol{a}})$.
@@
    Since $\exp(tX)=tX$ for $G=\mathbb{R}^{n+1}$, in coordinates we have $\xi_{X}=X^{0}\partial_{\bar{S}}+\sum_{i=1}^{n}X^{i}\partial_{\bar{a}_{i}}$. The connection must satisfy
@@
The principal $\mathbb{R}^{n+1}$-bundle structure $(\mathcal{M},\mathcal{B},\Xi,\mathbb{R}^{n+1})$ reveals that the redundancy in thermodynamic labeling: multiple values of $(\bar{S},\bar{\boldsymbol{a}})$ corresponding to the same physical quantum state $\sigma\in\mathcal{B}$ at fixed intensive parameters $\bar{\boldsymbol{\lambda}}$ is a gauge symmetry.
@@
Let $\Xi:\mathcal{M}\to\mathcal{B}$ be the principal quantum thermodynamic bundle. A global gauge transformation is a bundle automorphism $f:\mathcal{M}\to\mathcal{M}$, i.e., a diffeomorphism satisfying:
@@
In local coordinates $(\bar{S},\bar{\boldsymbol{a}},\bar{\boldsymbol{\lambda}})$ on $\mathcal{M}$, a gauge transformation $f\in\mathcal{G}(\mathcal{M})$ acts as
$f(\bar{S},\bar{\boldsymbol{a}},\bar{\boldsymbol{\lambda}})=\bigl(\bar{S}+\phi_{\bar{S}}(\bar{\boldsymbol{\lambda}}),\bar{\boldsymbol{a}}+\boldsymbol{\phi}_{\bar{\boldsymbol{a}}}(\bar{\boldsymbol{\lambda}}),\bar{\boldsymbol{\lambda}}\bigr),$ (84)
where $\phi_{\bar{S}}:\mathcal{B}\to\mathbb{R}$ and $\boldsymbol{\phi}_{\bar{\boldsymbol{a}}}:\mathcal{B}\to\mathbb{R}^{n}$ are smooth functions. This translates thermodynamic labels within each fiber while preserving the physical state: $\Xi(f(p))=\Xi(p)=\sigma$; and the intensive parameters: $\bar{\boldsymbol{\lambda}}(f(p))=\bar{\boldsymbol{\lambda}}(p)$.
Assuming the injectivity of the Gibbs map $\boldsymbol{\lambda}\mapsto\rho_{\boldsymbol{\lambda}}$ (see Remark 8), the equilibrium submanifold $\mathcal{E}$ intersects each fiber $F_{\sigma}$ at exactly one point. Equilibrium states therefore provide a natural gauge-fixing condition: for every physical quantum state $\sigma\in\mathcal{B}$ there exists a unique thermodynamic representative on $\mathcal{E}$, thereby eliminating the gauge redundancy carried by the fibers.
@@
Note that the horizontal distribution can be defined as the kernel of $\omega$:
@@
XI.2 Curvature on the principal bundle
In a local trivialization $\Xi^{-1}(U)\cong U\times\mathbb{R}^{n+1}$ of the principal $\mathbb{R}^{n+1}$-bundle $(\mathcal{M},\mathcal{B},\Xi,\mathbb{R}^{n+1})$, let $U\subset\mathcal{B}$ be an open set with coordinates $\{\bar{\lambda}_{k}\}_{k=1}^{n}$ (so $b_{k}=\bar{\lambda}_{k}$), and let the fiber $\mathbb{R}^{n+1}$ have coordinates $\{v_{\alpha}\}_{\alpha=0}^{n}$, where $v_{0}=\bar{S},$ and $v_{i}=\bar{a}_{i},\ \text{for }i=1,\dots,n$. The left-invariant Maurer–Cartan form $\theta$ on the abelian Lie group $G=(\mathbb{R}^{n+1},+)$ is